OpenStudy (anonymous):

How do I find a side, when only given 1 angle and 1 side? in an non-right angled triangle

OpenStudy (istim):

Are the angles and sides opposite each other?

OpenStudy (istim):

If so, use sin law

OpenStudy (istim):

Whereas sinA/a=sinB/b

OpenStudy (anonymous):



OpenStudy (anonymous):

@IsTim , I put a picture up

OpenStudy (istim):

is the angle at c?

OpenStudy (anonymous):

Yes,

OpenStudy (alekos):

you have an iscoseles triangle

OpenStudy (istim):

Well, one way you could figure this out is Sin Law. Look it up on purplemath, Khan's Academy or wolframalpha if you're new to the concept.

OpenStudy (anonymous):

@alekos, yes

OpenStudy (istim):

Do you understand, @Sulle26 ?

OpenStudy (alekos):

has it clicked yet?

OpenStudy (anonymous):

@IsTim, not really.... Can i find the other two angles by using 180-33.10 and then divide it by 2 to get those to angles, since it's an iscoceles triangle? @alekos

OpenStudy (istim):

You're looking for side lengths though, aren't you?

OpenStudy (istim):

But yes, Sulle26, you can do that to find the angle, which leads into using Sin Law.

OpenStudy (anonymous):

@IsTim, yes. So if i have at least 2 angles, I can find the remaining side, can't i ?

OpenStudy (istim):

Using Sin Law, yes.

OpenStudy (anonymous):

@IsTim, thanks!

OpenStudy (alekos):

you're heading in the right direction and remember the two unknown sides are also equal

OpenStudy (anonymous):

dont find side as simple as that

OpenStudy (anonymous):

@alekos Yup! I'll remember!

OpenStudy (istim):

@thenerdisback What do you mean?

OpenStudy (istim):

@Sulle26 I think this question is done. You can close it now.

OpenStudy (anonymous):

You can also use the Law of Cosines \[ c^2=a^2-2 a b \cos (c)+b^2=-2 a^2 \cos (c)+a^2+a^2=2 a^2 (1-\cos (c))\\ a^2=\frac{c^2}{2 (1-\cos (c))}=\frac{6.71^2}{2 (1-\cos (33.1 {}^{\circ}))}=138.72\\ a=\sqrt{138.72}=138.72 \]

OpenStudy (anonymous):

How is sqroot of 138.72 the same? but 11.77, right ? :-) @eliassaab

OpenStudy (istim):

Yes.

OpenStudy (anonymous):

I just used sin, and I got the same answer. Thanks everyone! :-)

OpenStudy (anonymous):

Sorry that was a misprint \[ c^2=a^2-2 a b \cos (c)+b^2=-2 a^2 \cos (c)+a^2+a^2=2 a^2 (1-\cos (c))\\ a^2=\frac{c^2}{2 (1-\cos (c))}=\frac{6.71^2}{2 (1-\cos (33.1 {}^{\circ}))}=138.72\\ a=\sqrt{138.72}=11.78 \]